-7x + y = -19
-2x + 3y = -19
Elimination method
Can you show work please. Having hard time. Thank you.

Found 3 solutions by ikleyn, Alan3354, macston:
Answer by ikleyn(52778)

.
We are here not for doing your work for you (or instead of you).

We are here to teach you.

And this technique was just shown to you in the preceding posts today.


Answer by Alan3354(69443)

-7x + y = -19
-2x + 3y = -19
Elimination method
===================
-21x +3y = -57 (1st eqn times 3)
-2x + 3y = -19
----------------------- Subtract
-19x = -38
x = 2
------------
-7x + y = -19
-14 + y = -19
y = -5

Answer by macston(5194)

.
-7x+y=-19 . Multiply by 3
-21x+3y=-57
-2x+3y=-19 . Subtract this from above (Eliminate 3y):
.
-19x=-38
x=2
.
Substitute x=2 in Equation 1:
-7x+y=-19
-7(2)+y=-19
-14+y=-19
y=5
.
CHECK (use Equation 2):
With x=2 and y=-5:
.
-2x+3y=-19
-2(2)+3(-5)=-19
-4+(-15)=-19
-19=-19
.